# A Unified Implementation of Quasi-Monte Carlo Generators, Randomization Routines, and Fast Kernel Methods

### v1

@@ Appendix A Proofs of theorems
Suppose $f\in H_{\alpha}$. Using the fact that (6) is a Mercer kernel and (Dick, 2009a, Theorem 14, Remark 19), we have
$\left\lVert f\right\rVert_{{\widetilde{H}}_{\alpha}}^{2}=\sum_{k\in{\mathbb{N}}_{0}}\left\lvert{\widehat{f}}(k)\right\rvert^{2}b^{\mu_{\alpha}(k)}\leq C_{\alpha}\sum_{k\in{\mathbb{N}}}b^{-\mu_{\alpha}(k)}<\infty$
for some $C_{\alpha}<\infty$, so $f\in{\widetilde{H}}_{K}$. ∎
Lemma A.1 (Walsh coefficients of low order monomials).
@@
${\widehat{f}}_{1}(k)$ =$=\begin{cases}1/2,&k=0\\ -2^{-a_{1}-2},&k=2^{a_{1}}\\ 0,&\mathrm{otherwise}\end{cases},$
${\widehat{f}}_{2}(k)$ =$=\begin{cases}1/3,&k=0\\ -2^{-a_{1}-2},&k=2^{a_{1}}\\ 2^{-a_{1}-a_{2}-3},&k=2^{a_{1}}+2^{a_{2}}\\ 0,&\mathrm{otherwise}\end{cases},$
${\widehat{f}}_{3}(k)$ =$=\begin{cases}1/4,&k=0\\ -2^{-a_{1}-2}+2^{-3a_{1}-5},&k=2^{a_{1}}\\ 3\cdot 2^{-a_{1}-a_{2}-4},&k=2^{a_{1}}+2^{a_{2}}\\ -3\cdot 2^{-a_{1}-a_{2}-a_{3}-5},&k=2^{a_{1}}+2^{a_{2}}+2^{a_{3}}\\ 0,&\mathrm{otherwise}\end{cases}.$
@@
The forms for ${\widehat{f}}_{1}$ and ${\widehat{f}}_{2}$ follow from (Dick and Pillichshammer, 2010, Example 14.2, Example 14.3). For $k=0$ and any $x\in[0,1)$ we have $\mathrm{wal}_{0}(x)=1$, so
@@
Assume $k\in{\mathbb{N}}$ going forward. For $k=2^{a_{1}}+k^{\prime}$ with $0\leq k^{\prime}<2^{a_{1}}$, (Fine, 1949, Equation 3.6) implies
@@
Using integration by parts and the fact that $J_{k}(0)=J_{k}(1)=0$
@@
${\widetilde{K}}_{\alpha}(x)=1+\sum_{1\leq\nu<\alpha}s_{\nu}(x)+{\widetilde{s}}_{\alpha}(x)$
where $s_{\nu}$ sums over all $k$ with #$\#k=\nu$ and ${\widetilde{s}}_{\alpha}$ sums over all $k$ with #$\#k\geq\alpha$. In (Baldeaux et al., 2012, Corollary 1) it was shown that
@@
from which ${\widetilde{K}}_{2}$ and ${\widetilde{K}}_{3}$ follow. We now find expressions for $s_{3}$ and ${\widetilde{s}}_{4}$ from which ${\widetilde{K}}_{4}$ follows.
Assume sums over $a_{i}$ are over ${\mathbb{N}}_{0}$ unless otherwise restricted. Recall from Lemma A.1 that
@@
$s^{[3}(x)=\sum_{a_{1}>a_{2}>a_{3}}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}+2^{a_{3}}}(x)}{2^{a_{1}+a_{2}+a_{3}+3}}=-\frac{4}{3}x^{3}+2x^{2}-\frac{2}{3}x+\frac{1}{24}\sum_{a_{1}}\frac{\mathrm{wal}_{2^{a_{1}}}(x)}{2^{3a_{1}}}.$
@@
$=\sum_{a_{1}>a_{2}>a_{3}>a_{4}}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}+2^{a_{3}}+2^{a_{4}}}(x)}{2^{a_{1}+a_{2}+a_{3}+a_{4}+4}}\sum_{0\leq k<2^{a_{4}}}\mathrm{wal}_{k}(x)$
If $x=0$ then
@@
$W_{7}-W_{6}$ $=\left(\frac{1}{2^{\beta(x)+1}}-\sum_{\beta(x)-1>a_{3}}\frac{1}{2^{a_{3}+2}}\right)\left(\sum_{a_{1}>a_{2}>\beta(x)-1}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}}(x)}{2^{a_{1}+a_{2}+2}}\right).$
@@
$-\frac{1}{48}\left(2-\beta(x)\right)\sum_{a_{1}}\frac{\mathrm{wal}_{2^{a_{1}}}(x)}{2^{3a_{1}}}-\frac{1}{294}\left(7\beta(x)+701t_{3}(x)\right)+\frac{5}{98}$

### v2

@@ Appendix B Proofs of theorems
Suppose $f\in H_{\alpha}^{\prime}$. Using (7) and [Dick, 2009a, Theorem 14, Remark 19], we have
$\left\lVert f\right\rVert_{H_{\alpha}}^{2}=\sum_{k\in{\mathbb{N}}_{0}}\left\lvert{\widehat{f}}(k)\right\rvert^{2}b^{\mu_{\alpha}(k)}\leq C_{\alpha}\sum_{k\in{\mathbb{N}}}b^{-\mu_{\alpha}(k)}<\infty$
for some $C_{\alpha}<\infty$, so $f\in H_{K}$. ∎
lemma B.1 (Walsh coefficients of low order monomials).
@@
${\widehat{f}}_{1}(k)$ $=\begin{cases}1/2,&k=0\\ -2^{-a_{1}-2},&k=2^{a_{1}}\\ 0,&\mathrm{otherwise}\end{cases},$
${\widehat{f}}_{2}(k)$ $=\begin{cases}1/3,&k=0\\ -2^{-a_{1}-2},&k=2^{a_{1}}\\ 2^{-a_{1}-a_{2}-3},&k=2^{a_{1}}+2^{a_{2}}\\ 0,&\mathrm{otherwise}\end{cases},$
${\widehat{f}}_{3}(k)$ $=\begin{cases}1/4,&k=0\\ -2^{-a_{1}-2}+2^{-3a_{1}-5},&k=2^{a_{1}}\\ 3\cdot 2^{-a_{1}-a_{2}-4},&k=2^{a_{1}}+2^{a_{2}}\\ -3\cdot 2^{-a_{1}-a_{2}-a_{3}-5},&k=2^{a_{1}}+2^{a_{2}}+2^{a_{3}}\\ 0,&\mathrm{otherwise}\end{cases}.$
@@
The forms for ${\widehat{f}}_{1}$ and ${\widehat{f}}_{2}$ follow from [Dick and Pillichshammer, 2010, Example 14.2, Example 14.3]. For $k=0$ and any $x\in[0,1)$ we have $\mathrm{wal}_{0}(x)=1$, so
@@
Assume $k\in{\mathbb{N}}$ going forward. For $k=2^{a_{1}}+k^{\prime}$ with $0\leq k^{\prime}<2^{a_{1}}$, [Fine, 1949, Equation 3.6] implies
@@
Using integration by parts and the fact that $J_{k}(0)=J_{k}(1)=0$,
@@
$K_{\alpha}(x)=\sum_{1\leq\nu<\alpha}s_{\nu}(x)+{\widetilde{s}}_{\alpha}(x)$
where $s_{\nu}$ sums over all $k$ with $k=\nu$ and ${\widetilde{s}}_{\alpha}$ sums over all $k$ with $k\geq\alpha$. In [Baldeaux et al., 2012, Corollary 1] it was shown that
@@
from which $K_{2}$ and $K_{3}$ follow. We now find expressions for $s_{3}$ and ${\widetilde{s}}_{4}$ from which $K_{4}$ follows.
Assume sums over $a_{i}$ are over ${\mathbb{N}}_{0}$ unless otherwise restricted. Lemma B.1 gives
@@
$s_{3}(x)=\sum_{a_{1}>a_{2}>a_{3}}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}+2^{a_{3}}}(x)}{2^{a_{1}+a_{2}+a_{3}+3}}=-\frac{4}{3}x^{3}+2x^{2}-\frac{2}{3}x+\frac{1}{24}\sum_{a_{1}}\frac{\mathrm{wal}_{2^{a_{1}}}(x)}{2^{3a_{1}}}.$
@@
$=\sum_{a_{1}>a_{2}>a_{3}>a_{4}}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}+2^{a_{3}}+2^{a_{4}}}(x)}{2^{a_{1}+a_{2}+a_{3}+a_{4}+4}}\sum_{0\leq k<2^{a_{4}}}\mathrm{wal}_{k}(x).$
If $x=0$, then
@@
$W_{7}-W_{6}$ $=\left(\frac{1}{2^{\beta(x)+1}}-\sum_{\beta(x)-1>a_{3}}\frac{1}{2^{a_{3}+2}}\right)\left(\sum_{a_{1}>a_{2}>\beta(x)-1}\frac{\mathrm{wal}_{2^{a_{1}}+2^{a_{2}}}(x)}{2^{a_{1}+a_{2}+2}}\right)$
@@
$-\frac{1}{48}\left(2-\beta(x)\right)\sum_{a_{1}}\frac{\mathrm{wal}_{2^{a_{1}}}(x)}{2^{3a_{1}}}-\frac{1}{294}\left(7\beta(x)+701t_{3}(x)\right)+\frac{5}{98},$